OpenStudy (brookemariexx):

Calculus! Please help f(x) = sqrt x+9; g(x) = 8x – 13 Find f(g(x)).

OpenStudy (igreen):

Doesn't look like calc to me.

OpenStudy (brookemariexx):

It's from my calculus class

OpenStudy (igreen):

Plug in 8x - 13 for 'x' in the first function. \(\sf \sqrt{8x - 13 + 9}\) Can you simplify it?

OpenStudy (brookemariexx):

Okay! like this? sqrt 8(8x-13)+9?

OpenStudy (brookemariexx):

sqrt 64x-95

OpenStudy (brookemariexx):

oh wait no

OpenStudy (brookemariexx):

2 sqrt 2x - 1

OpenStudy (igreen):

No, combine like terms first

OpenStudy (igreen):

-13 + 9 = ?

OpenStudy (brookemariexx):

-4

OpenStudy (igreen):

Yes..that leaves us with: \(\sf \sqrt{8x - 4}\)

OpenStudy (brookemariexx):

Okay thank you so much!

OpenStudy (igreen):

Yw